OpenStudy (jmansedg):

What is the slope of the line? y = (5/3)x - 5 A. -5 B. -5/3 C. 5/3 D. 5

OpenStudy (jmansedg):

a?

OpenStudy (luigi0210):

Well, your equation is already in \(\large y=mx+b \) (slope-intercept form), just look for what \(\large m \) is \(\Large y=\color{red}{\frac{5}{3}}x-5 \) \(\Large y=\color{red}{m}x+b \) :)

OpenStudy (jmansedg):

c?